# Mirror Speculative Decoding: Breaking the Serial Barrier in LLM Inference

### v1

@@ F.1 Target and Draft Sharding
For the experiments in Section 4, both target and draft models were distributed across eight Apple M2 Ultra systems  (Apple Inc., 2023a), each integrating a high-throughput GPU and a dedicated Neural Engine (NPU). We allocate the target to GPUs using Megatron-style tensor parallelism and the draft to NPUs using SPD-style sharding (see Section 3.3). Each M2 Ultra consists of a dual-die package connected internally by UltraFusion, a die-to-die interconnect providing up to 2.5 TB/s of bandwidth while presenting the system as a single logical GPU/NPU pair (Apple Inc., 2023a). Across machines, we organize the 8 nodes into groups of 2, linked by Thunderbolt 5 interconnects (up to 120 Gbps peak bandwidth) (Apple Inc., 2024). Groups are further connected through a high-speed network fabric, providing sufficient bandwidth for inter-group synchronization with sub-millisecond latency.
@@ F.2 Draft Model Configuration
The draft used in our experiments is a 0.6B-parameter model trained with the SPD architecture (Kim et al., 2025). It is organized into 16 transformer layers, divided into two contiguous segments of 8 layers each. Within every segment we instantiate $G_{D}{=}8$ parallel tracks, where track $g\in\{1,\dots,G_{D}\}$ is pinned to NPU $g$ and advances through its resident shard of the segment. Each track operates with a hidden size of 256 per shard. As in Section 3.3, there is no inter-NPU traffic within a segment. Synchronization occurs only twice per forward pass: once at the segment boundary to re-align tensor partitions, and once at the output to assemble logits for both main and lookahead streams.

### v2

@@ F.1 Target and Draft Sharding
For the experiments in Section 4, both target and draft models were distributed across eight Apple M2 Ultra systems  (apple_m2_ultra), each integrating a high-throughput GPU and a dedicated Neural Engine (NPU). We allocate the target to GPUs using Megatron-style tensor parallelism and the draft to NPUs using SPD-style sharding (see Section 3.3). Each M2 Ultra consists of a dual-die package connected internally by UltraFusion, a die-to-die interconnect providing up to 2.5 TB/s of bandwidth while presenting the system as a single logical GPU/NPU pair (apple_m2_ultra). Across machines, we organize the 8 nodes into groups of 2, linked by Thunderbolt 5 interconnects (up to 120 Gbps peak bandwidth) (apple_thunderbolt5). Groups are further connected through a high-speed network fabric, providing sufficient bandwidth for inter-group synchronization with sub-millisecond latency.
@@ F.2 Draft Model Configuration
The draft used in our experiments is a 0.6B-parameter model trained with the SPD architecture (Kim2025SPDSD). It is organized into 16 transformer layers, divided into two contiguous segments of 8 layers each. Within every segment we instantiate $G_{D}{=}8$ parallel tracks, where track $g\in\{1,\dots,G_{D}\}$ is pinned to NPU $g$ and advances through its resident shard of the segment. Each track operates with a hidden size of 256 per shard. As in Section 3.3, there is no inter-NPU traffic within a segment. Synchronization occurs only twice per forward pass: once at the segment boundary to re-align tensor partitions, and once at the output to assemble logits for both main and lookahead streams.